OpenStudy (anonymous):

5x/2 – 1/b2 = x/2b solve for x and simplify need to see work

OpenStudy (anonymous):

is it b ^2 nd after equal the 2b ???

OpenStudy (anonymous):

if so then 5x/2-1/b^2=x/2b or, 5x/2-x/2b=1/b^2 or, x/2b(5b-1)=1/b^2 or, x=2/{b(5b-1)}

OpenStudy (anonymous):

first 1

OpenStudy (anonymous):

so the 2nd one is ur ans..

OpenStudy (anonymous):

ok